find the equation of the line that is perpendicular to 9x-3y=-2 and contains the point (0,4). i know the answer is y=-1/3x+4 but i don't know how to get it.
Answer by mananth(16946)

9 x -3 y = -2
Find the slope of this line

-3 y = -9 x -2
Divide by -3
y = 3 x + 0.67
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 3

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be - 1/3

m= - 1/3 ,point ( 0 , 4 )
Find b by plugging the values of m & the point in
y=mx+b
4 = 0.00 + b
b= 4
m= - 1/3
The required equation is y =- 1/ 3x+4
m.ananth@hotmail.ca